# Equivariant Filter for Tightly Coupled LiDAR-Inertial Odometry*

### v1

@@ VI Experiments
In this section, we conduct a series of experiments to evaluate the accuracy and robustness of Eq-LIO. In all experiments, we compare Eq-LIO with the IEKF-based LIO algorithm and the EKF-based FAST-LIO2 algorithm. Furthermore, for fairness, all algorithms use the same tuning parameters in the experiments.
@@ VI-A Public Datasets
colspec=*4X[c,m], width=hline1,Z=1pt, hline2, cell2-Z1=font= Error(m) & FAST-LIO2 IEKF EqF
@@
colspec=X[1.2,c,m]*3X[c,m], width=hline1,Z=1pt, hline2, cell2-Z1=font= Error(m) & FAST-LIO2 IEKF EqF
@@ VI-B Private Datasets
1) Indoor Experiment: Robustness is a crucial characteristic of filter-based LiDAR-inertial odometry, referring to the system’s ability to withstand adverse external disturbances such as imperfect tuning parameters, degenerate motion scenes, or sustained intense movement. To induce significant changes in both rotation and velocity, the data collectors held the sensor in their hands and ran irregularly indoors, generating large and unpredictable movements. The data were collected using a Livox Avia LiDAR and its built-in MEMS IMU, with sampling frequencies of 100 Hz for LiDAR and 10 Hz for the IMU. Fig. 6 illustrates the attitude estimation during the experiment, revealing dramatic changes in a short period, making it ideal for examining the robustness of the proposed system. The results, depicted in Fig. 7, demonstrate that Eq-LIO produces the clearest mapping result and handles wall corners more effectively than IEKF. The mapping results of EKF-based FAST-LIO2 exhibit significant drift, while Eq-LIO shows less distortion on the walls, with the upper right corner of the map appearing clearer. This suggests that when handling highly nonlinear motion scenarios, the model based on Eq-LIO exhibits a superior linearization effect, making it more resilient to significant parameter errors and demonstrating enhanced robustness.
@@ VII Conclusion
This article presents Eq-LIO. To address the inconsistency and linearization point issues of traditional filtering algorithms, this work develops an equivariant filter for LIO systems, derives a new error definition based on the symmetry of the semi-direct product group, and incorporates gravity constraints. The algorithm achieves improved accuracy and robustness without increasing computational resources. Experiments on various public and private datasets show that, compared to existing methods, Eq-LIO can achieve higher accuracy and more importantly, robustness to challenging motion scenes at comparable speeds.

### v2

@@ VI Experiments
In this section, we conduct a series of experiments to evaluate the accuracy and robustness of Eq-LIO. In all experiments, we compare Eq-LIO with the IEKF-based LIO algorithm and the EKF-based FAST-LIO2 algorithm. Furthermore, for fairness, all algorithms use the same tuning parameters in the experiments. The transformation of the parameters between EKF and EqF is provided in the supplementary material[22].
@@ VI-A Public Datasets
colspec=*4X[c,m], width=hline1,Z=1pt, hline2, cell2-Z1=font= Error(m) & FAST-LIO2 IEKF Eq-LIO
@@
colspec=X[1.2,c,m]*3X[c,m], width=hline1,Z=1pt, hline2, cell2-Z1=font= Error(m) & FAST-LIO2 IEKF Eq-LIO
@@ VI-B Private Datasets
1) Indoor Experiment: Robustness is a crucial characteristic of filter-based LiDAR-inertial odometry, referring to the system’s ability to withstand adverse external disturbances such as imperfect tuning parameters, degenerate motion scenes, or sustained intense movement. To induce significant changes in both rotation and velocity, the data collectors held the sensor in their hands and ran irregularly indoors, generating large and unpredictable movements. The data were collected using a Livox Avia LiDAR and its built-in MEMS IMU, with sampling frequencies of 10 Hz for LiDAR and 100 Hz for the IMU. Fig. 6 illustrates the attitude estimation during the experiment, revealing dramatic changes in a short period, making it ideal for examining the robustness of the proposed system. The results, depicted in Fig. 7, demonstrate that Eq-LIO produces the clearest mapping result and handles wall corners more effectively than IEKF. The mapping results of EKF-based FAST-LIO2 exhibit significant drift, while Eq-LIO shows less distortion on the walls, with the upper right corner of the map appearing clearer. This suggests that when handling highly nonlinear motion scenarios, the model based on Eq-LIO exhibits a superior linearization effect, making it more resilient to significant parameter errors and demonstrating enhanced robustness.
@@ VII Conclusion
This article presents Eq-LIO. To address the inconsistency and linearization point issues of traditional filtering algorithms, this work develops an equivariant filter for LIO systems, derives a new error definition based on the symmetry of the semi-direct product group, and incorporates gravity constraints. The algorithm achieves improved accuracy and robustness without increasing computational resources. Experiments on various public and private datasets show that, compared to existing methods, Eq-LIO can achieve higher accuracy and more importantly, robustness to challenging motion scenes at comparable speeds. In future, the gravity state is supposed to be enbraced into the equivariant system.